# The Critique of Critique

### v1

@@ Abstract
Critique, as a natural language description for assessing the quality of model-generated content, has been proven to play an essential role in the training, evaluation, and refinement of Large Language Models (LLMs). However, there is a lack of principled understanding in evaluating the quality of the critique itself. In this paper, we pioneer the critique of critique, termed MetaCritique, which is a framework to evaluate the critique from two aspects, i.e., factuality as precision score and comprehensiveness as recall score. We calculate the harmonic mean of precision and recall as the overall rating called F1 score. To obtain a reliable evaluation outcome, we propose Atomic Information Units (AIUs), which describe the critique in a more fine-grained manner. MetaCritique takes each AIU into account and aggregates each AIU’s judgment for the overall score. Moreover, given the evaluation process involves intricate reasoning, our MetaCritique provides a natural language rationale to support each judgment. We construct a meta-evaluation dataset containing 300 critiques (2653 AIUs) across four tasks (question answering, reasoning, entailment, and summarization), and we conduct a comparative study to demonstrate the feasibility and effectiveness. Experiments also show superior critique judged by MetaCritique leads to better refinement, indicating generative artificial intelligence indeed has the potential to be significantly advanced with our MetaCritique. We will release relevant code and meta-evaluation datasets at https://github.com/GAIR-NLP/MetaCritique.
@@ 1 Introduction
Natural language critique has assumed a crucial role in advancing the development of Large Language Models (LLMs), ranging from the training of a more helpful and harmless model (Bai et al., 2022; OpenAI, 2023; Scheurer et al., 2023; Wu et al., 2023), alignment evaluation of model generations (Wang et al., 2023b; Zheng et al., 2023; Chan et al., 2023; Li et al., 2023) to the refinement of defective model outputs (Madaan et al., 2023; Gou et al., 2023; Ye et al., 2023; Akyurek et al., 2023). While a bunch of recent works are being done using generated critiques to assist in the development of large language models (Cui et al., 2023; Kim et al., 2023; Wang et al., 2023a; Li et al., 2023; Ke et al., 2023), there has not been enough emphasis on how to automatically and efficiently evaluate the quality of these critiques due to the following challenges: (i) quantification: establishing specific criteria to determine the critique rating, (ii) reliability: ensuring transparency to calculate the comparable score, and (iii) intricacy: grasping the complex relations among multiple concepts in the critique evaluation.
@@
In this paper, we pioneer the critique of critique, termed MetaCritique, to get over the above hurdles. An example of our MetaCritique is shown in Figure 1. Firstly, our MetaCritique tackles the quantification issue by establishing the criteria for critique evaluation, i.e., a meaningful critique should provide factual statements and comprehensive assessments, which will be quantified by two metrics we developed: precision and recall. Precision serves to gauge the accuracy of the critique’s content, ensuring each point is factual, while recall measures the extent to which the critique fully covers the necessary breadth of information, reflecting its comprehensiveness. Secondly, our MetaCritique addresses the reliability concern by introducing Atomic Information Units (AIUs). AIUs symbolize the fundamental segments of informative critique that can not be divided further. Our MetaCritique converts the critique-level evaluation into AIU-level evaluation to minimize ambiguity in the evaluation process. Subsequently, MetaCritique aggregates these AIU-level outcomes to produce the overall score, thereby guaranteeing the transparency of the scoring process. Lastly, inspired by how complex reasoning problems are alleviated using LLMs (Zhou et al., 2023), we attempt to resolve the intricacy problem by generating a natural language rationale step-by-step, i.e., Chain-of-Thought (CoT) (Wei et al., 2022), for each AIU-level judgment. As a result, it enhances the reliability of judgment and facilitates human involvement in the evaluation loop.
@@
To validate the feasibility and effectiveness of our MetaCritique, we create a meta-evaluation dataset covering four tasks (question answering, reasoning, entailment, and summarization) across 16 public datasets, and we conduct a comparative analysis of human-written and LLM-generated critiques. Our MetaCritique achieved near-human performance so MetaCritique can serve as an effective critique evaluation for human annotation and LLM’s reflexion. Besides, MetaCritique can identify high-quality critiques, which lead to improved results via iterative refinement. It indicates that MetaCritique can enhance the alignment of existing LLMs.
Our contributions are summarized as follows:
We are the pioneers in prioritizing critique evaluation and introducing the critique of critique, termed MetaCritique, involving the principles of precision and recall.
Our MetaCritique is more interpretable and transparent, wherein critiques are decomposed into AIUs, and natural language rationale is provided for each judgment.
We collect a meta-evaluation dataset involving question answering, reasoning, entailment, and summarization to confirm the feasibility and effectiveness of our MetaCritique, which have been publicly released.
@@ 2 MetaCritique
MetaCritique aims to evaluate the quality of hypothesis critique by generating the critique of hypothesis critique. It involves three steps, i.e., reference generation, AIU extraction, and hypothesis critiquing as illustrated in Figure 2. To facilitate the detailed description of MetaCritique, we first introduce some important concepts.
Question denotes a user’s query or instruction that prompts the generative model to produce a pertinent and insightful response.
Model-Generated Answer describes the textual content created by the generative model as a reaction to the question.
Reference Answer is a ground-truth answer for the question. It is used for the precision score.
Hypothesis Critique is the natural language feedback to point out errors of the model-generated answer and provide actionable suggestions. It can be written by either human annotators or LLMs. It will be evaluated in this paper.
Reference Critique represents an ideal critique to comment accurately and thoroughly on the model-generated answers. It is used to calculate the recall score.
@@
Atomic Information Unit (AIU) is the smallest unit that can self-sufficiently convey a piece of information without any additional context. MetaCritique decomposes a critique into AIUs to reduce the ambiguity of the evaluation process and improve the transparency of the outcome. Generating a numeric score for the critique depends on the information volume it encompasses. For example, a precision score considers the volume of correct information in the critique. However, information, being intangible, renders the task of determining the volume of information unfeasible (Porat, 1977; Soofi, 1994). By extracting and counting AIUs from a critique, we can approximate the volume of information as the number of AIUs because AIUs are the most fundamental elements.
2.1 Reference Generation
@@
Reference is essential for most text generation evaluation (Papineni et al., 2002; Lin, 2004; Zhang* et al., 2020; Yuan et al., 2021). Also, it is crucial for our MetaCritique. We need reference answers to calculate the precision score, and reference critiques to calculate the recall score. However, reference answers or reference critiques are always unavailable because they require significant human effort. To address this issue, we adopt the GPT-4 generated content as a proxy of the references as in previous works (OpenAI, 2023; Zheng et al., 2023; Peng et al., 2023; Cui et al., 2023; Li et al., 2023).
Considering that GPT-4 can make mistakes with improper instructions, we refined our instruction via manual review, whereby we tested various instructions on a series of examples and assessed the alignment between the generations with human preference. Given appropriate prompts, GPT-4 can achieve remarkable results (6.51/7 for reference answer and 6.56/7 for reference critique under the human evaluation in Table 3). We provide detailed prompting instructions in Table 7.
2.2 AIU Extraction
@@
AIU extraction aims to split a critique into AIUs. It is similar to some prior works such as verifiable claims extraction for factuality detection (Chern et al., 2023) and atomic content units extraction for summarization evaluation (Liu et al., 2023b). It has been proven that suitably prompted GPT-4 can precisely conduct such a task as mentioned in Chern et al. (2023). Inspired by their success, we implement AIUs extraction by prompting GPT-4. We provide detailed instructions in Table 9.
2.3 Hypothesis Critiquing
@@
We prompt strong LLMs with strong instruction-following capability to implement this evaluation. We follow the idea of CoT reasoning to design the instructions for prompting LLMs. Firstly, the LLM needs to find the necessary information to verify the AIU from the question, model-generated answer, and reference answer. Then, the LLM attempts to explain and reason about whether the AIU is factual or not. Finally, the LLM states the conclusion. Detailed prompting instructions with demonstrations are in Table 10.
After checking each AIU in the hypothesis critique, we aggregate the precision score as the proportion of factual AIUs relative to the total count of AIUs in the hypothesis critique as follows:
where is the number of all AIUs extracted from the hypothesis critique, and is the -th AIU of the hypothesis critique.
@@
We prompt strong LLMs with strong instruction-following capability to implement this evaluation. We follow the idea of CoT reasoning to design the instructions for prompting LLMs. Firstly, the LLM attempts to explain and reason about whether the AIU from the reference critique is mentioned or logically inferred from the hypothesis critique. Then, the LLM states the conclusion whether the AIU is entailed or not. Detailed prompting instructions with demonstrations are in Table 11.
After checking each AIU in the reference critique, we denote the recall score of the hypothesis critique as the ratio of entailed AIUs to all AIUs from the reference critique as follows:
where is the number of all AIUs extracted from the reference critique and is the -th AIU of the reference critique.
@@
We introduce the F1 Score as an overall assessment score of hypothesis critique to harmonize the precision and recall. It is calculated as the harmonic mean of the precision score and recall score as follows:
@@ 3 Meta-Evaluation Dataset
In this section, we elaborate on constructing a meta-evaluation dataset for comparing MetaCritique with its counterparts, along with presenting its statistical features.
@@ 3.1 Collection of Question and Model-Generated Answer
To get broad coverage of NLP domains, we collect ready-made question-answer pairs from the Shepherd (Wang et al., 2023a) dataset, which consists of question-answer-critique triads for generating critique across various domains. The model-generated answer in this dataset has reasonable errors, so that the critique can be generated to improve the response. We carefully extract some data as our meta-evaluation data. As a result, we collect data from four domains: entailment, reasoning, question answering, and summarization, across 16 datasets: Entailment Bank (Dalvi et al., 2021), e-SNLI (Camburu et al., 2018), Adversarial NLI (Nie et al., 2020), ECQA (Aggarwal et al., 2021), CosmosQA (Huang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), RACE (Lai et al., 2017), SIQA (Sap et al., 2019), TriviaQA (Joshi et al., 2017), Natural Question (Kwiatkowski et al., 2019), BoolQ (Clark et al., 2019), GPT-3 summarization (Goyal et al., 2022), DeFacto (Liu et al., 2023a).
@@ 3.2 Collection of Hypothesis Critique
Hypothesis critique is the evaluation object in this paper. It can be written by human annotators (abbreviated to Hypo.h in tables) and LLMs (abbreviated to Hypo.l in tables). It is valuable to evaluate human-written critique, as it can help to understand the common shortcomings of human annotations. Also, it is necessary to understand how well the LLM-generated critiques perform because a large number of LLM-generated critiques have been used for training, evaluation, and refinement.
For human-written critique, we collect the ready-made human-written critique from Shepherd dataset. There is only one human-written critique for a question-answer pair. For LLM-generated critiques, we collect two critiques for each question-answer pair. These two critiques will be further used to conduct pairwise comparisons. We do not conduct the pairwise comparison between human-written critiques and LLM-generated critiques to avoid potentially misleading or unfair outcomes that could arise from inherent differences in their writing styles. In detail, we randomly select two different LLMs from WizardLM (Xu et al., 2023) (13B and 70B), GPT-3.5, LLaMA-2 (Touvron et al., 2023) (chat-13B and chat-70B), Vicuna (Chiang et al., 2023) (13B and 33B), and SelFee (Ye et al., 2023) (13B). Then, we respectively prompt the selected LLM to generate the critique. We use the same prompt as the reference critique generation (Table 7) for GPT-3.5. As for other LLMs, detailed prompting instructions can be found in Table 8.
@@ 3.3 Collection of Human Annotation
We collect human labeling results for precision and recall tasks driven by two objectives: (1) to evaluate the performance of different LLMs to execute the precision and recall tasks. (2) to meta-evaluate MetaCritique and its counterparts in the light of human judgments.
Before annotation, we prepare the reference answer, reference critique and extracted AIUs via GPT-4 as introduced in Section 2.1 and 2.2. We engaged two postgraduate students for performing the precision and recall task. As illustrated in Figure 2, they replace the GPT-4 to provide solely binary labels without explanation. A third postgraduate meticulously reviews the work of the first two annotators, ensuring accuracy and resolving any discrepancies. This structured process is essential for maintaining the reliability of our research data. Lastly, we can calculate the precision score, recall score, and F1 score via these annotated results. These scores, representing human judgments, are named as gold scores in this paper.
@@ 3.4 Statistics
Table 1 shows the statistics of our meta-evaluation dataset. The number of AIUs can be viewed as the information volume of the critique. We find that human-written critiques contain less information than LLM-generated critiques (including reference critiques), nearly half of LLM-generated critiques.
| AIU Level (Micro) | |||
| Meta-P | Meta-R | Meta-F1 | |
| Hypo.h | 87.61 | 48.72 | 62.62 |
| Hypo.l | 71.85 | 53.28 | 61.19 |
| Critique Level (Macro) | |||
| Meta-P | Meta-R | Meta-F1 | |
| Hypo.h | 85.37 | 50.97 | 58.24 |
| Hypo.l | 71.07 | 54.37 | 58.20 |
According to human annotation, we calculate the gold scores of precision, recall, and F1 score at the AIU level and critique level. The results are shown in Table 2. We find that human-written critiques get higher precision scores yet lower recall scores than LLM-generated critiques. It indicates that LLMs are even more prone to creating fictional information, while human is hardly possible to make such mistakes. Also, it is worth noting that human misses more key information than LLMs. This might be because human-written critiques produce fewer AIUs than LLM-generated critiques.

### v2

@@ Abstract
Critique, as a natural language description for assessing the quality of model-generated content, has played a vital role in the training, evaluation, and refinement of LLMs. However, a systematic method to evaluate the quality of critique is lacking. In this paper, we pioneer the critique of critique, termed MetaCritique, which builds specific quantification criteria. To achieve a reliable evaluation outcome, we propose Atomic Information Units (AIUs), which describe the critique in a more fine-grained manner. MetaCritique aggregates each AIU’s judgment for the overall score. Moreover, MetaCritique delivers a natural language rationale for the intricate reasoning within each judgment. Lastly, we construct a meta-evaluation dataset covering 4 tasks across 16 public datasets involving human-written and LLM-generated critiques. Experiments demonstrate that MetaCritique can achieve near-human performance. Our study can facilitate future research in LLM critiques based on our following observations and released resources: (1) superior critiques judged by MetaCritique can lead to better refinements, indicating that it can potentially enhance the alignment of existing LLMs; (2) the leaderboard of critique models reveals that open-source critique models commonly suffer from factuality issues; (3) relevant code and data are publicly available at https://github.com/GAIR-NLP/MetaCritique to support deeper exploration; (4) an API at PyPI with the usage documentation in Appendix C allows users to assess the critique conveniently.
@@ 1 Introduction
Natural language critique has assumed a crucial role in advancing the development of Large Language Models (LLMs), ranging from the training of a more helpful and harmless model (Bai et al., 2022; OpenAI, 2023; Scheurer et al., 2023; Wu et al., 2023), alignment evaluation of model generations (Wang et al., 2023b; Zheng et al., 2023; Chan et al., 2023; Li et al., 2023) to the refinement of defective model outputs (Madaan et al., 2023; Gou et al., 2023; Ye et al., 2023; Akyurek et al., 2023). While a bunch of recent works are being done using generated critiques to assist in the development of LLMs (Cui et al., 2023; Kim et al., 2023; Wang et al., 2023a; Li et al., 2023; Ke et al., 2023), there has not been enough emphasis on how to automatically and efficiently evaluate the quality of these critiques due to the following challenges: (i) quantification: establishing specific criteria to qualify the critique rating, (ii) reliability: ensuring transparency to calculate the comparable score, and (iii) intricacy: grasping the complex relations among multiple concepts in the critique evaluation.
@@
In this paper, we pioneer the critique of critique, termed MetaCritique, to get over the above hurdles. An example of MetaCritique is shown in Figure 1. Firstly, MetaCritique tackles the quantification issue by establishing specific criteria, i.e., a meaningful critique should provide factual statements and comprehensive assessments, which will be quantified by two metrics: precision and recall. Precision serves to gauge the accuracy of the critique’s content, ensuring each point is factual, while recall measures the extent to which the critique fully covers the necessary breadth of information, reflecting its comprehensiveness. Secondly, MetaCritique addresses the reliability concern by introducing Atomic Information Units (AIUs). AIUs symbolize the fundamental segments of informative critique that cannot be divided further. MetaCritique converts the critique-level evaluation into AIU-level evaluation to minimize ambiguity in the evaluation process. Subsequently, MetaCritique aggregates these AIU-level outcomes to produce the overall score, thereby guaranteeing the transparency of the scoring process. Lastly, inspired by the fact that complex reasoning problems can be alleviated by using LLMs with Chain-of-Thought (CoT) (Wei et al., 2022), we attempt to resolve the intricacy problem by generating a natural language rationale step by step for each AIU-level judgment. As a result, it can enhance the reliability of judgment and facilitate human involvement in the evaluation loop.
@@
Given the absence of a dataset for critique evaluation, we curate a meta-evaluation dataset covering 4 tasks (question answering, reasoning, entailment, and summarization) across 16 public datasets, involving human-written and LLM-generated critiques. Our MetaCritique achieves near-human performance, indicating that MetaCritique can help understand human annotation and LLM’s reflection. Besides, MetaCritique can identify high-quality critiques, which lead to improved results via iterative refinement. This indicates that MetaCritique can enhance the alignment of existing LLMs. A leaderboard of critique models also aids in identifying the pros and cons of various critique models. In conclusion, MetaCritique can potentially advance the progress of LLMs.
@@ 2 MetaCritique
MetaCritique evaluates the quality of a hypothesis critique by generating its critique, i.e., the critique of critique. It involves three steps: (1) reference generation, (2) AIU extraction, and (3) hypothesis critiquing as illustrated in Figure 2.
@@
To facilitate the detailed description of MetaCritique, we first introduce some important concepts. Question denotes a user’s query or instruction that prompts LLMs to produce a pertinent and insightful response. Model-generated Answer describes the textual content created by LLMs as a reaction to the question. Reference Answer is a ground-truth answer for the question. Hypothesis Critique is the natural language feedback to point out errors of the model-generated answer and provide actionable suggestions. It can be written by either human annotators or LLMs. Reference Critique is an ideal critique to comment accurately and thoroughly on the model-generated answers. Atomic Information Unit (AIU) is the smallest unit that can self-sufficiently convey a piece of information. It can help reduce the ambiguity of the evaluation process and improve the transparency of the evaluation outcome. Generating a numeric score for the critique depends on the information volume it encompasses. For example, a precision score considers the volume of correct information in the critique. However, information, being intangible, renders the task of determining the volume of information unfeasible (Porat, 1977; Soofi, 1994). By extracting and counting AIUs from a critique, we can approximate the volume of information as the number of AIUs because AIUs are the most fundamental elements of critiques.
2.1 Step 1: Reference Generation
@@
Reference is essential for most text generation evaluation (Papineni et al., 2002; Lin, 2004; Zhang* et al., 2020; Yuan et al., 2021). Also, it is crucial for our MetaCritique. We need reference answers to calculate the precision score, and reference critiques to calculate the recall score. However, reference answers or critiques are always unavailable because they require significant human effort. To address this issue, we adopt the GPT-4 generated content as a proxy of the references as in previous works (OpenAI, 2023; Zheng et al., 2023; Peng et al., 2023; Cui et al., 2023; Li et al., 2023). Moreover, we conduct human evaluations to ensure the quality of the GPT-4 generated content. We provide detailed prompting instructions in Table 8.
2.2 Step 2: AIU Extraction
@@
AIU extraction aims to split a critique into AIUs. It is similar to some prior works such as verifiable claims extraction for factuality detection (Chern et al., 2023) and atomic content units extraction for summarization evaluation (Liu et al., 2023b). It has been proven that suitably prompted GPT-4 can precisely conduct such a task as mentioned in Chern et al. (2023). Inspired by their success, we implement AIUs extraction by prompting GPT-4. Besides, we conduct human evaluations to verify the quality of the GPT-4 extraction. We provide detailed prompting instructions in Table 10.
2.3 Step 3: Hypothesis Critiquing
@@
We prompt LLMs with strong instruction-following capability to implement this evaluation. We follow the idea of CoT reasoning to design the instructions. Firstly, the LLM needs to find the necessary information to verify the AIU. Then, the LLM explains and reasons whether the AIU is factual or not. Finally, the LLM states the conclusion. Detailed prompting instructions with demonstrations are shown in Table 11.
After checking each AIU in the hypothesis critique, we denote the precision score as the proportion of factual AIUs relative to the total count of AIUs in the hypothesis critique.
@@
We prompt LLMs guided by the CoT reasoning to perform this evaluation. Firstly, the LLM analyses if the AIU from the reference critique is mentioned or logically inferred from the hypothesis critique. Subsequently, the LLM states whether the AIU is entailed or not. Detailed prompting instructions with demonstrations are in Table 12.
After checking each AIU in the reference critique, we denote the recall score of the hypothesis critique as the ratio of entailed AIUs to all AIUs from the reference critique.
@@
We introduce the F1 Score as an overall assessment score, which harmonizes the precision score and recall score as follows:
@@ 3 Meta-Evaluation Dataset
In this section, we elaborate on constructing a meta-evaluation dataset and human annotation for evaluating the critique evaluation, along with presenting its statistical features. We first collect questions, model-generated answers containing flaws, and human-written or LLM-generated critiques (hypothesis critique). Subsequently, we use GPT-4 to generate critiques and answers as references and extract AIUs from critiques. Finally, human annotators use these data to complete precision and recall tasks for each AIUs as shown in Figure 2.
@@ 3.1 Collection of Question and Model-Generated Answer
To get broad coverage of NLP domains, we collect ready-made question-answer pairs from the Shepherd (Wang et al., 2023a) dataset, which consists of question-answer-critique triads for generating critique across various domains. The model-generated answer in this dataset has reasonable errors, so that the critique can be generated to improve the response. We carefully extract some data to exclude the tasks that need specific tools to find errors, like code generation. As a result, we collect data from four domains: entailment, reasoning, question answering, and summarization, across 16 datasets: Entailment Bank (Dalvi et al., 2021), e-SNLI (Camburu et al., 2018), Adversarial NLI (Nie et al., 2020), ECQA (Aggarwal et al., 2021), CosmosQA (Huang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), RACE (Lai et al., 2017), SIQA (Sap et al., 2019), TriviaQA (Joshi et al., 2017), Natural Question (Kwiatkowski et al., 2019), BoolQ (Clark et al., 2019), GPT-3 summarization (Goyal et al., 2022), DeFacto (Liu et al., 2023a).
@@ 3.2 Collection of Hypothesis Critique
Hypothesis critique can be written by human annotators (abbreviated to Hypo.h in tables) and LLMs (abbreviated to Hypo.l in tables). It is valuable to evaluate human-written critique, as it can help to understand the common shortcomings of human annotations. Also, it is necessary to understand how well the LLM-generated critiques perform because LLM-generated critiques have been widely used for training, evaluation, and refinement.
For human-written critique, we use one ready-made critique from the Shepherd dataset. For LLM-generated critiques, we generate two critiques for each question-answer pair. These two critiques will be further used to conduct pairwise comparisons. We do not conduct the pairwise comparison between human-written critiques and LLM-generated critiques to avoid potentially misleading outcomes that could arise from their writing styles. In detail, we randomly select two LLMs from WizardLM (Xu et al., 2023) (13B and 70B), GPT-3.5, LLaMA-2 (Touvron et al., 2023) (chat-13B and chat-70B), Vicuna (Chiang et al., 2023) (13B and 33B), and SelFee (Ye et al., 2023) (13B). Then, we respectively prompt the selected LLM to generate one critique. We use the same prompt as the reference critique generation (Table 8) for GPT-3.5. As for other LLMs, the prompt is shown in Table 9.
@@ 3.3 Collection of Human Annotation
We collect human labeling results for precision and recall tasks driven by two objectives: (1) to evaluate the performance of different LLMs to execute the precision and recall tasks (Exp. II). (2) to meta-evaluate MetaCritique and its counterparts in the light of human judgments (Exp. III-IV).
Before annotation, we prepare the reference answer, reference critique and extracted AIUs via GPT-4 as introduced in Section 2.1 and 2.2. We engaged two postgraduate students to perform the precision and recall task. As shown in Figure 2, they replace the GPT-4 to provide solely binary labels without explanation. A third postgraduate meticulously reviews the work of the first two annotators, ensuring accuracy and resolving any discrepancies. This process is essential for maintaining the reliability of our research data. Lastly, we can calculate the MetaCritique scores via these annotated results. These scores (shown in Table 6) represent human judgments (named as gold scores).
@@ 3.4 Statistics
Table 1 shows the statistics of our meta-evaluation dataset. We collect 100 question-answer pairs. Each pair has 1 human-written critique, 2 LLM-generated critiques, and 1 reference critique. We find that the number of AIUs in human-written critiques is less than half of it in LLM-generated critiques or reference critiques. This implies that human-written critiques possibly contain less information than LLM-generated critiques.